# Christian Döding [^2] Institut für Numerische Simulation, Universität Bonn, Friedrich-Hirzebruch-Allee 7, D-53115 Bonn, Germany    Barbara Verfürth [^2] Institut für Numerische Simulation, Universität Bonn, Friedrich-Hirzebruch-Allee 7, D-53115 Bonn, Germany

### v2

@@ 1 Introduction
Metamaterials are artificially structured media with spatially and/or temporally varying material properties that enable wave phenomena beyond those found in natural materials, such as amplification, negative refraction indices, non-reciprocity, and cloaking [4, 5, 6, 9, 10, 11, 14, 25]. In recent years, temporal and spatiotemporal metamaterials have attracted significant attention due to their intriguing physical behaviour and potential applications in optics, acoustics, and elasticity. The mathematical analysis of such materials has become an active area of research, ranging from Morgenthaler’s early work on time-varying media [22] to more recent studies in time-harmonic regimes [12, 15, 16, 27, 32, 33] and in the context of homogenization theory [18, 23, 24, 28, 29, 31].
@@
This work is devoted to the higher-order homogenization of temporally modulated electromagnetic metamaterials characterized by a spatially homogeneous but temporally heterogeneous permittivity. Electromagnetic wave propagation in such media is governed by Maxwell’s equations, which give rise to wave-type equations with time-dependent coefficients for the electric and magnetic fields. Using a formal two-scale asymptotic expansion, we derive effective wave equations for both fields. In addition to the leading-order homogenized equations, we carry out higher-order homogenization by deriving governing equations for the first- and second-order correctors of the effective solution. These higher-order terms yield refined approximations of the effective description and capture nonlocal physical effects as light propagates through the medium. We observe that the homogenized equations differ for the electric and magnetic fields due to the distinct placement of the time-dependent coefficients in their respective wave equations. Furthermore, a key finding of this study is the derivation of initial conditions that require careful treatment and introduce temporal interface conditions for the homogenized equations. Recasting our results within the Maxwell framework leads to both local and nonlocal effective material laws. Nonlocality means that the corresponding constitutive relation involves derivatives of the electric field, aligning with the literature, c.f. [17, 21, 24]. While our primary focus is electromagnetics, the generality of our approach extends naturally to other wave systems in temporally modulated media, such as acoustics and elasticity.
@@
Our analysis generalizes and unifies results from previous studies. For instance, the work [24] addresses a case corresponding to the electric field setting considered here and derives second-order effective equations in the time-frequency domain, which are consistent with our results in physical space. Our contribution complements theirs by providing a general derivation of the homogenized equations, adding initial conditions, and extending the analysis to the magnetic field case. Higher-order spatial homogenization for wave-type equations has been studied in [13, 30]. However, these settings do not cover the time-dependent case considered in our study, since the coefficients vary in space rather than in time. Although a formal interchange of space and time leads to similar homogenized equations, this approach does not cover initial values and temporal interface conditions or the case of space dimension larger than one. In the context of spatiotemporal metamaterials, higher-order homogenization has been explored in the works [28, 29], which study material modulations along traveling wave trajectories. Roughly speaking, our setting belongs to case of infinite modulation velocity in [28, 29] which is excluded therein.
@@
At leading order, an effective description for purely time-dependent dielectric composites was derived in [23], where the authors obtain the harmonic mean as the effective permittivity for a two-component system with piecewise constant modulations. Their setting also corresponds to our electric field case. Although our framework assumes smooth modulations, it can be applied to the composite setting of [23]. In contrast, [31] examines time-varying acoustic media and finds that the arithmetic mean is the effective coefficient. At first, this seems to contradict [23], but in [31], the authors consider time-varying compressibility. Via the derivation of the wave equation, this corresponds to a different placement of the time-varying coefficient, which leads to a different homogenized coefficient, as highlighted in our analysis.
@@ 3 Homogenization via asymptotic expansion
$\mathbf{u}_{j}(x,t,\tau)=\overline{\mathbf{u}}_{j}(x,t)+\widetilde{\mathbf{u}}_{j}(x,t,\tau),\quad j\in\mathbb{N}_{0}.$
@@
for all $(x,t)\in\mathbb{R}^{d}\times\mathbb{R}_{+}$. In the subsequent sections we first consider the equations for the electric displacement (2.8) (Section 4), which we frequently refer to as the electric case. Then we tackle the system (2.9) describing the propagation of the magnetic field (Section 5) which is referred to the magnetic case.
Clearly, for (3.1) to hold we need to assume that the solution $\mathbf{u}_{\eta}$ of (2.9) and (2.8), respectively, depends smoothly on $\eta$. This can be expected as long as the permittivity coefficient $\varepsilon_{\eta}$ and the initial data $\mathbf{v}^{\mathrm{e}}_{0}$, $\mathbf{v}^{\mathrm{e}}_{1}$ and $\mathbf{v}^{\mathrm{m}}_{0}$, $\mathbf{v}^{\mathrm{m}}_{1}$ are sufficiently smooth. However, in this work we do not aim at a rigorous justification of the formal asymptotic expansion (3.1), but at the derivation and mathematical underpinning of the effective local and nonlocal material laws as electromagnetic waves propagate through time-varying metamaterials. In this sense, our results are purely formal, but several classical homogenization techniques can be used to rigorously justify our findings, e.g., the method of two-scale convergence [1] or other related homogenization methods [2, 7, 8, 19, 26]. The techniques used therein have mainly been applied to spatial multiscale problems, but a suitable exchange of space and time might allow to treat time-dependent problems in a similar way. In fact, the media oscillate rapidly with respect to a one-dimensional variable, namely time, and homogenization in one dimension is an extensively studied area. Finally, for an application of these techniques, it is crucial that our media are constant in space and that the initial values do not contain any spatial multiscale features. This is expressed by the small spatial modulation frequencies imposed by the initial values in (2.8) and (2.9), so that no coupling with the fast temporal scale can occur.
@@ 4.1 Effective electric behavior
Integration over $(0,1)$ w.r.t. $\tau$ and using the periodicity of $\mathbf{u}_{2}$ yields the compatibility condition
@@ 4.2 Electric first order correction
We can calculate $\partial_{\tau}\mathbf{u}_{2}(x,0,0)$ explicitly. For this, we recall the $\mathcal{O}(\eta)$-term of the asymptotic expansion in the PDE, see (4.3). Integrating with respect to $\tau$, we obtain
@@ 4.3 Electric higher order corrections
$\partial_{\tau\tau}\mathbf{u}_{4}+2\partial_{t\tau}\mathbf{u}_{3}+\partial_{tt}\mathbf{u}_{2}-\varepsilon^{-1}\Delta\mathbf{u}_{2}=0.$
From there we obtain the compatibility condition
$0=\int_{0}^{1}\partial_{tt}\mathbf{u}_{2}-\varepsilon^{-1}\Delta\mathbf{u}_{2}\,d\tau=\int_{0}^{1}\partial_{tt}\overline{\mathbf{u}}_{2}-\varepsilon^{-1}\Delta\overline{\mathbf{u}}_{2}-\varepsilon^{-1}\theta\Delta\partial_{tt}\mathbf{u}_{0}\,d\tau.$
Replacing $\Delta\partial_{tt}\mathbf{u}_{0}$ by $\varepsilon_{\hom}^{-1}\Delta^{2}\mathbf{u}_{0}$ due to (4.6), this yields the homogenized equation with a (nonlocal) correction
@@ 4.4 Pure macroscopic description of the electric field
$\varepsilon_{\hom}\partial_{tt}\overline{\mathbf{u}}^{(2)}-\Delta\overline{\mathbf{u}}^{(2)}-\eta^{2}\varepsilon_{\mathrm{cor}}\Delta^{2}\overline{\mathbf{u}}^{(2)}=-\eta^{3}\varepsilon_{\mathrm{cor}}\Delta^{2}\overline{\mathbf{u}}_{1}-\eta^{4}\varepsilon_{\mathrm{cor}}\Delta^{2}\overline{\mathbf{u}}_{2}.$
Neglecting the $\mathcal{O}(\eta^{3})$-terms on the right hand side we can define $\mathbf{u}^{(2)}$ as the solution of the initial value problem
@@ 4.5 The electric field $\mathbf{E}$
Figure 4.1: The fields $\mathbf{E}$ (left) and $\mathbf{D}$ (right). Model setting as described in Section 7.
@@ 5 The magnetic case
Next, we homogenize the magnetic field and consider the initial value problem (2.9), where the electric permittivity occurs between the two time derivatives in the wave equation. Pretending that there is only one spatial dimension, we can exchange space and time, so that this case is similar to the homogenization of the (one-dimensional) wave equation with spatially multiscale coefficients (in divergence form), and we expect to obtain similar results. We follow the same strategy as in the electric case, plugging the asymptotic expansion into the PDE and initial conditions (2.9) and comparing powers in $\eta$. However, due to the different type of wave equations, we also get different problems for the effective solution and the higher order corrections compared to the electric case.
@@ 5.1 Effective magnetic behavior
$\partial_{\tau}(\varepsilon\partial_{t}\mathbf{u}_{0})+\partial_{\tau}(\varepsilon\partial_{\tau}\mathbf{u}_{1})=0.$
In contrast to the electric case, this does not imply $\mathbf{u}_{1}$ to be independent of $\tau$. Rather, we calculate (due to linearity) $\mathbf{u}_{1}$ in dependence on $\mathbf{u}_{0}$ and obtain
$\mathbf{u}_{1}(x,t,\tau)=\chi(\tau)\partial_{t}\mathbf{u}_{0}(x,t)+\overline{\mathbf{u}}_{1}(x,t),$
where $\chi$ is the unique $1$-periodic solution of the cell-problem
@@
Integrating the equation from 0 to $1$ w.r.t. $\tau$ leads to the compatibility condition
$\int_{0}^{1}\partial_{t}(\varepsilon\partial_{t}\mathbf{u}_{0})+\partial_{t}(\varepsilon\partial_{\tau}\mathbf{u}_{1})-\Delta\mathbf{u}_{0}\,d\tau=0,$
where we used the periodicity of $\varepsilon$$\mathbf{u}_{1}$, and $\mathbf{u}_{2}$. Since $\mathbf{u}_{0}$ is independent of $\tau$ and $\partial_{\tau}\mathbf{u}_{1}=(\partial_{t}\mathbf{u}_{0})\partial_{\tau}\chi$, we deduce
@@ 5.2 Magnetic first order correction
where $\xi$ is the unique $1$-periodic solution of the second cell-problem
@@
$\partial_{\tau}(\varepsilon\partial_{\tau}\mathbf{u}_{3})+\partial_{t}(\varepsilon\partial_{\tau}\mathbf{u}_{2})+\partial_{\tau}(\varepsilon\partial_{t}\mathbf{u}_{2})+\partial_{t}(\varepsilon\partial_{t}\mathbf{u}_{1})-\Delta\mathbf{u}_{1}=0.$
Integration over $(0,1)$ w.r.t. $\tau$ gives the compatibility condition
@@
$0=\partial_{t}\mathbf{u}_{1}(x,0,\tau)+\partial_{\tau}\mathbf{u}_{2}(x,0,\tau).$
Integration over $(0,1)$ w.r.t. $\tau$ yields due to the periodicity of $\mathbf{u}_{2}$

### v3

@@ 1 Introduction
Metamaterials are artificially structured media with spatially and/or temporally varying material properties that enable wave phenomena beyond those found in natural materials, such as amplification, negative refraction indices, non-reciprocity, and cloaking. In recent years, temporal and spatiotemporal metamaterials have attracted significant attention due to their intriguing physical behavior and potential applications in optics, acoustics, and elasticity. We refer to the seminal works [3, 6, 7, 8, 11, 13, 14, 18, 25, 30] for a comprehensive introduction and overview of the theoretical foundations, design, and applications of spatiotemporal metamaterials from both physical and mathematical perspectives. The mathematical analysis of temporal metamaterials, in particular, has become an active area of research in recent decades. This field ranges from Morgenthaler’s early work on time-varying media [27] to more recent studies in time-harmonic regimes [2, 16, 19, 20, 32, 37, 38]. The present work contributes to this research field from the standpoint of mathematical homogenization theory. As we elaborate below, it is closely related to studies on spatiotemporal metamaterials [12, 17, 22, 28, 29, 33, 34, 35, 36], in which different but conceptually related models of modulated media have been analyzed. Before discussing these connections, we first outline our model setting, describe the homogenization approach we exploit, and summarize our main contributions to the theory of homogenization for temporal metamaterials.
@@
In this work we consider higher-order homogenization of temporally modulated electromagnetic metamaterials characterized by a spatially homogeneous but temporally heterogeneous permittivity. Electromagnetic wave propagation in such media is governed by Maxwell’s equations, which give rise to wave-type equations with time-dependent coefficients for the electric and magnetic fields. Using a formal two-scale asymptotic expansion, we derive effective wave equations for both fields. In addition to the leading-order homogenized equations, we carry out higher-order homogenization by deriving governing equations for the first- and second-order correctors of the effective solution. These higher-order terms yield refined approximations of the effective description and capture nonlocal physical effects as light propagates through the medium. We observe that the homogenized equations differ for the electric and magnetic fields due to the distinct placement of the time-dependent coefficients in their respective wave equations. Furthermore, a key finding of this study is the derivation of initial conditions that require careful treatment and introduce temporal interface conditions for the homogenized equations. Recasting our results within the Maxwell framework leads to both local and nonlocal effective material laws. Nonlocality means that the corresponding constitutive relation involves derivatives of the electric field, aligning with the literature, c.f. [21, 26, 29]. While our primary focus is electromagnetics, the generality of our approach extends naturally to other wave systems in temporally modulated media, such as acoustics and elasticity.
@@
Next, we discuss how our analysis generalizes and unifies results from previous studies, contributing to the existing literature. The work [29] addresses a case corresponding to the electric field setting considered here and derives second-order effective equations in the time-frequency domain, which are consistent with our results in physical space. Our contribution complements theirs by providing a general derivation of the homogenized equations, adding initial conditions, and extending the analysis to the magnetic field case. Higher-order spatial homogenization for wave-type equations has been studied in [17, 35]. However, these settings do not cover the time-dependent case considered in our study, since the coefficients vary in space rather than in time. Although a formal interchange of space and time leads to similar homogenized equations, this approach does not cover initial values and temporal interface conditions or the case of space dimension larger than one. Spatiotemporal metamaterials with modulations along traveling-wave trajectories were studied at the effective level in [22], and higher-order homogenization was subsequently investigated in [33, 34]. Roughly speaking, our setting belongs to case of infinite modulation velocity in [33, 34] which is excluded therein.
@@
At leading order, an effective description for purely time-dependent dielectric composites was derived in [28], where the authors obtain the harmonic mean as the effective permittivity for a two-component system with piecewise constant modulations. Their setting also corresponds to our electric field case. Although our framework assumes smooth modulations, it can be applied to the composite setting of [28]. In contrast, [36] examines time-varying acoustic media and finds that the arithmetic mean is the effective coefficient. At first, this seems to contradict [28], but in [36], the authors consider time-varying compressibility. Via the derivation of the wave equation, this corresponds to a different placement of the time-varying coefficient, which leads to a different homogenized coefficient, as highlighted in our analysis. Finally, we note that spatiotemporal metamaterials were recently studied in [12] using a space-harmonic approach, which in certain cases allows for the derivation of exact solutions to the corresponding Maxwell system.
@@ 2 Problem derivation
We emphasize that assuming $\varepsilon_{\eta}$ to be scalar and real corresponds to an idealized model of an isotropic, lossless metamaterial, since a non-zero imaginary part would account for material absorption. While allowing $\varepsilon_{\eta}$ to be complex-valued does not alter our analytical approach, and the resulting homogenized equations remain of the same form, the anisotropic case is much more involved and not covered by our approach.
@@ 3 Homogenization via asymptotic expansion
$\mathbf{u}_{j}(x,t,\tau)=\overline{\mathbf{u}}_{j}(x,t)+\widetilde{\mathbf{u}}_{j}(x,t,\tau),\quad j\in\mathbb{N}_{0}.$ (3.2)
@@
for all $(x,t)\in\mathbb{R}^{d}\times\mathbb{R}_{+}$. The decomposition (3.2) into macroscopic and microscopic components is motivated by separating the dominant, energy-carrying wave dynamics from the rapid, fine-scale oscillations induced by temporal modulation. Due to the zero-mean condition (3.3) these fine-scale contributions average out over each period. In the subsequent sections we first consider the equations for the electric displacement (2.8) (Section 4), which we frequently refer to as the electric case. Then we tackle the system (2.9) describing the propagation of the magnetic field (Section 5) which is referred to the magnetic case.
Clearly, for (3.1) to hold we need to assume that the solution $\mathbf{u}_{\eta}$ of (2.9) and (2.8), respectively, depends smoothly on $\eta$. This can be expected as long as the permittivity coefficient $\varepsilon_{\eta}$ and the initial data $\mathbf{v}^{\mathrm{e}}_{0}$, $\mathbf{v}^{\mathrm{e}}_{1}$ and $\mathbf{v}^{\mathrm{m}}_{0}$, $\mathbf{v}^{\mathrm{m}}_{1}$ are sufficiently smooth. Beyond assuming the formal expansion (3.1), let us briefly discuss the regularity requirements on the permittivity and initial conditions. It suffices to assume that the permittivity is bounded, i.e., $\varepsilon_{\eta}\in L^{\infty}(0,T;\mathbb{R}_{+})$. Under this assumption, the homogenized coefficients are well-defined, and the subsequent formal derivations are justified. In the magnetic case, weak formulations in time are required to rigorously handle the equations. For the initial values $\mathbf{v}^{\mathrm{e}}_{0},\mathbf{v}^{\mathrm{e}}_{1}$ and $\mathbf{v}^{\mathrm{m}}_{0},\mathbf{v}^{\mathrm{m}}_{1}$ we assume regularity in Sobolev spaces $H^{k}(\Omega)$ with $k$ large enough to guarantee well-posedness of the subsequent homogenized wave equations in $L^{2}(0,T,H^{1}_{0}(\Omega))$, c.f. [15]
However, in this work we do not aim at a rigorous justification of the formal asymptotic expansion (3.1), but at the derivation and mathematical underpinning of the effective local and nonlocal material laws as electromagnetic waves propagate through time-varying metamaterials. In this sense, our results are purely formal, but several classical homogenization techniques can be used to rigorously justify our findings, e.g., the method of two-scale convergence [1] or other related homogenization methods [4, 9, 10, 23, 31]. The techniques used therein have mainly been applied to spatial multiscale problems, but a suitable exchange of space and time might allow to treat time-dependent problems in a similar way. In fact, the media oscillate rapidly with respect to a one-dimensional variable, namely time, and homogenization in one dimension is an extensively studied area. Finally, for an application of these techniques, it is crucial that our media are constant in space and that the initial values do not contain any spatial multiscale features. This is expressed by the small spatial modulation frequencies imposed by the initial values in (2.8) and (2.9), so that no coupling with the fast temporal scale can occur.
@@ 4.1 Effective electric behavior
Integration over $(0,1)$ w.r.t. $\tau$ and using the periodicity of $\mathbf{u}_{2}$ yields
@@ 4.2 Electric first order correction
We can calculate $\partial_{\tau}\mathbf{u}_{2}(x,0,0)$ explicitly. For this, we recall the $\mathcal{O}(\eta)$-term of the asymptotic expansion in the PDE, see (4.3), and integrate it with respect to $\tau$ to obtain
@@ 4.3 Electric higher order corrections
$\partial_{\tau\tau}\mathbf{u}_{4}+2\partial_{t\tau}\mathbf{u}_{3}+\partial_{tt}\mathbf{u}_{2}-\varepsilon^{-1}\Delta\mathbf{u}_{2}=0.$ (4.13)
Integrating (4.13) over $(0,1)$ w.r.t. $\tau$ and using the periodicity of $\mathbf{u}_{3}$ and $\mathbf{u}_{4}$ yields
$0=\int_{0}^{1}\partial_{tt}\mathbf{u}_{2}-\varepsilon^{-1}\Delta\mathbf{u}_{2}\,d\tau=\int_{0}^{1}\partial_{tt}\overline{\mathbf{u}}_{2}-\varepsilon^{-1}\Delta\overline{\mathbf{u}}_{2}-\varepsilon^{-1}\theta\Delta\partial_{tt}\mathbf{u}_{0}\,d\tau.$ (4.14)
Replacing $\Delta\partial_{tt}\mathbf{u}_{0}$ by $\varepsilon_{\hom}^{-1}\Delta^{2}\mathbf{u}_{0}$ in (4.14) due to (4.6), this yields the homogenized equation with a (nonlocal) correction
@@ 4.4 Pure macroscopic description of the electric field
$\varepsilon_{\hom}\partial_{tt}\overline{\mathbf{u}}^{(2)}-\Delta\overline{\mathbf{u}}^{(2)}-\eta^{2}\varepsilon_{\mathrm{cor}}\Delta^{2}\overline{\mathbf{u}}^{(2)}=-\eta^{3}\varepsilon_{\mathrm{cor}}\Delta^{2}\overline{\mathbf{u}}_{1}-\eta^{4}\varepsilon_{\mathrm{cor}}\Delta^{2}\overline{\mathbf{u}}_{2}.$ (4.20)
Neglecting the $\mathcal{O}(\eta^{3})$-terms on the right hand side of (4.20) we can define $\mathbf{u}^{(2)}$ as the solution of the initial value problem
@@ 4.5 The electric field $\mathbf{E}$
Figure 4.1: Full-wave simulation of the electric fields $\mathbf{E}$ (left) and $\mathbf{D}$ (right) with permittivity $\varepsilon_{\eta}(t)=\big(2+\sin(2\pi t/\eta)\big)^{-1}$, $\eta=0.025$ and initial values $\mathbf{v}^{\mathrm{e}}_{0}=w(x,0)$, $\mathbf{v}^{\mathrm{e}}_{1}=\partial_{t}w(x,0)$ where $w(x,t)=e^{-200(t-x)^{2}}\cos(\omega_{0}(t-x))$, $\omega_{0}=0.01$.
@@ 5 The magnetic case
Next, we homogenize the magnetic field and consider the initial value problem (2.9), where the electric permittivity occurs between the two time derivatives in the wave equation. Pretending that there is only one spatial dimension, we can exchange space and time, so that this case is similar to the homogenization of the (one-dimensional) wave equation with spatially multiscale coefficients (in divergence form), and we expect to obtain similar results. We follow the same strategy as in the electric case, plugging the asymptotic expansion into the PDE and initial conditions (2.9) and comparing powers in $\eta$. However, due to the different type of wave equations, we also get different problems for the effective solution and the higher order corrections compared to the electric case. In addition, three distinct cell problems on the unit interval $(0,1)$ arise describing the microscopic behavior of the first- and second-order correctors. To facilitate the reading of the subsequent derivation, we collect the cell problems and their respective roles in Table 1.
solution cell-problem role/determines
$\chi$ $\partial_{\tau}(\varepsilon\partial_{\tau}\chi)=-\partial_{\tau}\varepsilon,\quad\displaystyle\int_{0}^{1}\chi\,d\tau=0$ microscopic components $\widetilde{\mathbf{u}}_{1}$ and $\widetilde{\mathbf{u}}_{2}$ interface conditions for $\overline{\mathbf{u}}_{1}$ and $\overline{\mathbf{u}}_{2}$
$\xi$ $\partial_{\tau}(\varepsilon\partial_{\tau}\xi)=-\partial_{\tau}(\varepsilon\chi),\quad\displaystyle\int_{0}^{1}\xi\,d\tau=0$ microscopic component $\widetilde{\mathbf{u}}_{2}$ interface condition for $\overline{\mathbf{u}}_{2}$ $\xi(0)=-\theta_{0}$ with $\theta_{0}$ from (4.12)
$\zeta$ $\partial_{\tau}(\varepsilon\partial_{\tau}\zeta)=-\partial_{\tau}(\varepsilon\xi)+\chi\varepsilon_{\hom},\quad\displaystyle\int_{0}^{1}\zeta\,d\tau=0$ only used as auxiliary problem
Table 1: Overview of the cell-problems arising in the derivation of the magnetic homogenized equations.
@@ 5.1 Effective magnetic behavior
$\partial_{\tau}(\varepsilon\partial_{t}\mathbf{u}_{0})+\partial_{\tau}(\varepsilon\partial_{\tau}\mathbf{u}_{1})=0.$ (5.1)
In contrast to the electric case, (5.1) does not imply $\mathbf{u}_{1}$ to be independent of $\tau$. Rather, we calculate (due to linearity) $\mathbf{u}_{1}$ in dependence on $\mathbf{u}_{0}$ and obtain
$\mathbf{u}_{1}(x,t,\tau)=\chi(\tau)\partial_{t}\mathbf{u}_{0}(x,t)+\overline{\mathbf{u}}_{1}(x,t).$
Here $\chi$ is the unique $1$-periodic solution of the first cell-problem, c.f. Table 1,
@@
Integrating (5.4) from 0 to $1$ w.r.t. $\tau$ leads to
$\int_{0}^{1}\partial_{t}(\varepsilon\partial_{t}\mathbf{u}_{0})+\partial_{t}(\varepsilon\partial_{\tau}\mathbf{u}_{1})-\Delta\mathbf{u}_{0}\,d\tau=0,$ (5.5)
where we used the periodicity of $\varepsilon$$\mathbf{u}_{1}$, and $\mathbf{u}_{2}$. Since $\mathbf{u}_{0}$ is independent of $\tau$ and $\partial_{\tau}\mathbf{u}_{1}=(\partial_{t}\mathbf{u}_{0})\partial_{\tau}\chi$, we deduce from (5.5)
@@ 5.2 Magnetic first order correction
where $\xi$ is the unique $1$-periodic solution of the second cell-problem, c.f. Table 1,
@@
$\partial_{\tau}(\varepsilon\partial_{\tau}\mathbf{u}_{3})+\partial_{t}(\varepsilon\partial_{\tau}\mathbf{u}_{2})+\partial_{\tau}(\varepsilon\partial_{t}\mathbf{u}_{2})+\partial_{t}(\varepsilon\partial_{t}\mathbf{u}_{1})-\Delta\mathbf{u}_{1}=0.$ (5.10)
Integrating (5.10) over $(0,1)$ w.r.t. $\tau$ gives
@@
$0=\partial_{t}\mathbf{u}_{1}(x,0,\tau)+\partial_{\tau}\mathbf{u}_{2}(x,0,\tau).$ (5.11)
Integration of (5.11) over $(0,1)$ w.r.t. $\tau$ yields due to the periodicity of $\mathbf{u}_{2}$